OpenStudy (anonymous):

3(x + 1) = −2(x − 1) − 4

OpenStudy (anonymous):

3x+3=-2x+2-4 3x=-2x-5 5x=-5 x=-1

OpenStudy (anonymous):

thank you!! One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? could you help me again please?

OpenStudy (anonymous):

Sure, just a second.

OpenStudy (anonymous):

576

OpenStudy (anonymous):

thank you :) i have some more questions if you could help Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

That one's just like the one I did earlier. I'll help you with it, but how bout you try it first?

OpenStudy (anonymous):

-2x+(-6) = -2x+(-1)-4

OpenStudy (anonymous):

Okay, almost. Look at that negative one. You forgot to distribute the 2 to it. You got the negative, but forgot the two. The rest is right, except for the -1. -2x+(-6) = -2x+(-2)-4 Kepp going :)